Solve (d^2y)/(dx^2)+y=sec^2x


1
Expert's answer
2021-07-20T09:29:30-0400

"{d^2y\over dx^2}+y=sec^2(x)"

But "sec^2(x)=tan^2(x)+1"

Rewriting the differential equation, we have

"{d^2y\over dx^2}+y=tan^2(x)+1......................(i)"

The corresponding homogenous linear equation is

"{d^2y\over dx^2}+y=0........................(ii)"

To solve this homogenous linear equation "(ii)" ,Find the roots of characteristic equation

"k^2+1=0 \implies" "k_1=-i" or "k_2=i"

The two roots of the characteristic equation are purely imaginary

"\therefore y(x)=C_1sin(x|k_1|)+C_2cos(x|k_2|)"

"\implies y(x)=C_1sin(x)+C_2cos(x)...................(iii)"

Equation "(iii)" is our general solution

Now we should solve the inhomogeneous equation "(i)" Using variation of parameters method

Suppose "C_1" is function of "x" and "C_2" is also a function of "x"

The general solution is:

"\implies y(x)=C_1(x)sin(x)+C_2(x)cos(x).................(iv)"

Where "C_1(x)" and "C_2(x)" by method of variation of parameters ,we find the solution from the system

"y_1(x){d\over dx}C_1(x)+y_2(x){d\over dx}C_2(x)=0"


"{d\over dx}C_1(x){d\over dx}y_1(x)+{d\over dx}C_2(x){d\over dx}y_2(x)=f(x)"


Where "y_1(x)" and "y_2(x)" are linearly independent solutions of linear Ordinary differential equations and "f(x)" is the free term and are given as

"y_1(x)=sin(x)" , "y_2(x)=cos(x)" and "f(x)=tan^2(x)+1"

Substituting to the system above, we have

"sin(x){d\over dx}C_1(x)+cos(x){d\over dx}C_2(x)=0"


"{d\over dx}C_1(x){d\over dx}sin(x)+{d\over dx}C_2(x){d\over dx}cos(x)=tan^2(x)+1"


But "{d\over dx}sin(x)=cos (x)" and "{d\over dx}cos(x)=-sin(x)"

Substituting to our system, we get:

"sin(x){d\over dx}C_1(x)+cos(x){d\over dx}C_2(x)=0"


"cos(x){d\over dx}C_1(x)-sin(x){d\over dx}C_2(x)=tan^2(x)+1"

Solving the system simultaneously, We get

"{d\over dx}C_1(x)={1\over cos(x)}" and "{d\over dx}C_2(x)=-{sin(x)\over cos^2(x)}"

let's solve "{d\over dx}C_1(x)={1\over cos(x)}" and "{d\over dx}C_2(x)=-{sin(x)\over cos^2(x)}"

"{d\over dx}C_1(x)={1\over cos(x)} \implies dC_1(x)={dx\over cos(x)}"

Integrating both sides we have

"\int dC_1(x)=\int {dx\over cos(x)}\implies C_1(x)=C_3-{log(sin(x)-1)\over 2}+{log(sin(x)+1)\over 2}"

"{d\over dx}C_2(x)=-{sin(x)\over cos^2(x)} \implies dC_2(x)=-{sin(x)\over cos^2(x)}dx"

Integrating both sides we have

"\int dC_2(x)=-\int{sin(x)\over cos^2(x)}dx\implies C_2(x)=C_4-{1\over cos(x)}"

Now, substitute "C_1(x)" and "C_2(x)" to the general solution "(iv)"

"y(x)=C_1(x)sin(x)+C_2(x)cos(x)"


"=(C_3-{log(sin(x)-1)\over 2}+{log(sin(x)+1)\over 2})sin(x)+(C_4-{1\over cos(x)})(x)cos(x)"


"\therefore y(x)=C_2cos(x)+(C_1-{log(sin(x)-1)\over 2}+{log(sin(x)+1)\over 2})sin(x)-1"